11=1/3a+2 solve and check
Answer by BedrockFist(9)

Isolate the variable. (Subtract 2 from both sides) 11-2=1/3a+2-2
Then you are left with 9 = 1/3a
Divide both sides be 1/3.
Then you get 27=a (Or a=27)
CHECK:
11=1/3(27)+2
11=9+2
11=11